Fermi National Laboratory
 

Operation Summary for March 31, 2010 - April 2, 2010

 

Wednesday March 31

The day shift began with Operations monitoring the Tevatron (TeV) store 7718 (169.88E30) and the Recycler stash (135.5E10), with the Antiproton Source stacking (19.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations established beam to the Meson MTest experiment at 11:45 AM.

During the evening shift, Operations terminated store 7718 at 9:40 PM and then put the TeV into HEP shot setup. Operations established store 7719 at 11 PM with an average initial luminosity (AIL) of 339.0E30.

Thursday April 1

The midnight shift began with Operations monitoring store 7719 (239.52E30) and the stash (45.6E10), with Pbar stacking (31.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that the MI extraction power supply (LAM52) tripped off at 1:15 AM. EE Support personnel came in and found a blown fuse inside the supply. Pbar resumed stacking at 3:35 AM.

During the day shift, Operations terminated store 7719 at 1:05 PM. TeV personnel began a series of machine studies.

During the evening shift, the TeV studies concluded and Operations put the TeV into shot setup at 6 PM. Operations established store 7722 at 7:15 PM with an AIL of 291.0E30.

Friday April 2

The midnight shift began with Operations monitoring store 7722 (131.78E30) and the stash (151.0E10), with Pbar stacking (18.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Booster extraction power supply (MP02) tripped off at 3:08 AM on a bad vacuum indication; no vacuum pumps had tripped off. It reset. The I- Source tripped off at 7:52 AM; it reset.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Recycler personnel to conduct some studies and FESS personnel to bring the main pond water pump back online at MI-10. Operations will terminate store 7722 around noon. Cryo system technicians will then begin work on various systems.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/02/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory